# ClassContrast: Bridging the Spatial and Contextual Gaps for Node Representations

### v1

@@ B.2 ClassContrast Homophily Matrices
B.3 Proof of the Theorem and New Homophily Metrics
In Section 4.4, we have defined homophily and discussed the relation between ClassContrast vectors and the homophily notion. Here we give details how to generalize this idea to give finer homophily notions for graphs. First, we give the proof of Theorem 4.7.
Theorem B.1.
For an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, let $\vec{\alpha}_{1}(v)$ be the spatial vector defined in Appendix B. Let $\widehat{\alpha}_{1}(v)$ be the vector where the entry corresponding to class of $v$ is set to 00. Then,
@@ B.3 Proof of the Theorem and New Homophily Metrics
[…]
@@
where the last equality follows by Equation 1. The proof follows. ∎
This perspective inspires different ways to generalize the homophily concept by using ClassContrast vectors. Notice that in ClassContrast matrices above, we employed a classwise grouping to measure class interactions. If we do not use any grouping for the nodes, we get natural generalizations of existing homophily ratios. First, we define higher homophily by using the ratio of the number of nodes in the $2$-neighborhood $\mathcal{N}_{2}(v)$ with the same class to $|\mathcal{N}_{2}(v)|$.
Definition B.2 (Higher Homophily).
Given $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with $\mathcal{C}:\mathcal{V}\to\{1,2,\dots,N\}$ representing node classes. Let $\eta_{2}(v)$ be the number of nodes in $\mathcal{N}_{2}(v)$ in the same class with $v$. Then, homophily ratio of $\mathcal{G}$ is defined as
@@
By applying similar ideas to the proof of Theorem 4.7, we obtain the following result.
Theorem B.3.
For an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, let $\vec{\alpha}_{2}(v)$ be the spatial vector defined above. Let $\widehat{\alpha}_{2}(v)$ be the vector where the entry corresponding to the class of $v$ is set to 00. Then,
@@
Definition B.4 (Contextual Homophily).
Given $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with $\mathcal{C}:\mathcal{V}\to\{1,2,\dots,N\}$ representing node classes. Let $\vec{\beta}(v)$ be the contextual vector as defined in Section 4.2. Let $\widehat{\beta}(v)$ be the vector where the entry corresponding to the class of $v$ is set to 00.Then, define the contextual homophily of $\mathcal{G}$ as
@@
Similar homophily metrics, as introduced in (Luan et al. 2023; Lim et al. 2021b; Zhu et al. 2020; Jin et al. 2022), were employed to investigate the concept of heterophily in graph representation learning. In particular, our ClassContrast features inherently encompass this information, enabling the ML classifier to adapt accordingly. This explains the outstanding performance of ClassContrast in both homophilic and heterophilic settings for node classification and link prediction.
Appendix C Generalizations of ClassContrast
C.1 Inductive and Transductive Settings
@@
To clarify the differences between the inductive and transductive settings in graph-based learning, consider a given graph $\mathcal{G}$, with datasets $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$. In the inductive setting, all test nodes and their connected edges are removed to create the training subgraph $\mathcal{G}_{\text{train}}$. The model is trained exclusively on $\mathcal{G}_{\text{train}}$ and only gains access to the complete graph $\mathcal{G}$ during the testing stage. Conversely, in the transductive learning approach, the model has access to the entire graph $\mathcal{G}$ during training; however, the labels of the test nodes remain hidden throughout this process. It is noteworthy that any dataset configured for transductive learning can be adapted to the inductive setting by excluding test nodes and their connecting edges during the training phase. However, converting from the inductive to the transductive setting is not generally feasible. For further details, see the references (Ciano et al. 2021; Arnold, Nallapati, and Cohen 2007).
C.2 ClassContrast for Transductive Setting and Iterated Predictions
So far, we outlined our ClassContrast vectors for simplicity in inductive setting. To adapt to the transductive setting, we make adjustments without altering contextual vectors. In transductive setting, test node labels are hidden during training, but connection information to training nodes is available. We introduce a new "unknown" class $\mathcal{C}_{u}$ for test nodes with unknown labels, considering them as neighbors to training nodes. Each node $u\in\mathcal{V}$ is represented by a $N+1$ dimensional vector $\vec{\alpha}^{0}_{1}(u)=[a^{0}_{10}\ \ a^{0}_{11}\ \ a^{0}_{12}\ \ \dots\ \ a^{0}_{1N}]$, where $a^{0}_{10}$ is the count of neighboring test nodes (unknown labels) of node $u$, and $a^{0}_{1j}$ is the count of neighboring training nodes in class $\mathcal{C}_{j}$ for $1\leq j\leq N$. Similar representations are defined for $\vec{\alpha}_{1i}(u)$, $\vec{\alpha}_{1o}(u)$, and $\vec{\alpha}_{2}(u)$. The superscript 00 indicates no iterations have occurred yet.
Iterated Predictions: Recall that the ultimate goal in the node classification problem is to predict the labels of new (test) nodes. After we obtain all ClassContrast vectors for training and test nodes above, we let our ML model make a prediction for each test node $v\in\mathcal{V}_{test}$. Let $\mathcal{P}_{0}:\mathcal{V}_{test}\to\{\mathcal{C}_{1},\dots,\mathcal{C}_{N}\}$ be our predictions. Hence, we have a label for each node in our graph $\mathcal{G}$.
@@ C.2 ClassContrast for Transductive Setting and Iterated Predictions
C.3 Spatial Embeddings on Directed, Weighted Graphs
@@
where $a^{i}_{kj}$ is the count of k-hop neighbors incoming to $u$ belonging to class $\mathcal{C}_{j}$ while $a^{o}_{kj}$ is the count of neighbors outgoing from $u$ belonging to class $\mathcal{C}_{j}$ (See Figure 2(a)).
@@
Table 11: The dimension of the ClassContrast embedding used for each dataset in our model
@@
CORA CITESEER PUBMED OGBN-ARXIV OGBN-MAG TEXAS CORNELL WISCONSIN CHAMELEON SQUIRREL
Dimension 42 36 112 120 1,524 30 30 30 20 20
Appendix D Details of ClassContrast Embeddings
@@
In all datasets, we basically used the same method to obtain our vectors, however, when the graph type (directed, undirected), or node attribute vector format varies, our methodology naturally adapts the corresponding setting as detailed below. Note that details of OGBN datasets can be found at (Hu et al. 2020) and at Stanford’s Open Graph Benchmark site  [^1] https://ogb.stanford.edu/docs/nodeprop/.
@@
CORA: The CORA dataset is a directed graph of scientific publications classified into one of the $7$ classes. Each node $u$ represents a paper and comes with a binary ($0/1$) vector $X_{u}$ of length 1433 indicating the presence/absence of the corresponding word in the paper from a dictionary of 1433 unique words. The task is to predict the subject area of the paper, e.g. Neural Networks, Theory, Case-Based. The directed graph approach is used to extract the attribute vector, $\gamma^{0}(u)$, from the CORA dataset. Recall that this is transductive setting, and the first row is an $8$ dimensional spatial vector $\alpha^{0}_{1i}(u)$ (Section 4.1) where the first $7$ entries represent the count of citing (incoming) papers from the corresponding class and the $8^{th}$ entry is the count of unknown citing paper $(v_{i}\in D_{test})$ (Section C.2). The second row $\alpha^{0}_{1o}(u)$ is defined similarly using the count of cited (outgoing) papers of the corresponding class. The third and fourth row of the spatial feature vector is obtained similarly using the second neighborhood information of each node for citing and cited paper respectively. For the first iteration $\gamma^{1}(u)$, the same setup is followed ignoring the $8^{th}$ entry of each row because there is no unknown class now.
@@
For the contextual vector $\vec{\beta}_{1}(u)$, we follow the landmark approach described in Section 4.2. We define the first landmark vector $\xi^{1}_{j}$ which is a binary word vector of length 1433 such that the entry is $1$ if the corresponding word is present in any of binary vectors $\mathcal{X}_{u}$ belonging to class $\mathcal{C}_{j}$, and the entry is 00 otherwise. Then, each entry $b_{1j}$ of $\vec{\beta}_{1}(u)$ is the count of common words between $\mathcal{X}_{u}$ and $\xi^{1}_{j}$ for each class, which produces a $7$ dimensional contextual vector. Similarly, for the contextual vector $\vec{\beta}_{2}(u)$, we use a more selective landmark vector $\xi^{2}_{j}$ is defined as a binary word vector of length 1433 indicating the presence/absence of the corresponding word in at least $10\%$ nodes in the class $\mathcal{C}_{j}$. Hence, the initial vector $\gamma^{0}(u)$ is $46$ dimensional ($32$ spatial, $14$ contextual), in the next iterations, $\gamma^{1}(u),\gamma^{2}(u)$ are both $42$-dimensional ($28$ spatial, $14$ contextual).
@@
OGBN-ARXIV: The OGBN-ARXIV dataset is a directed graph, representing the citation network between all Computer Science (CS) arXiv papers indexed by MAG (Wang et al. 2020). Each node is an arXiv paper and each directed edge indicates that one paper cites another one. Each paper comes with a 128-dimensional vector obtained by averaging the embeddings of words in its title and abstract. The embeddings of individual words are computed by running the skip-gram model (Mikolov et al. 2013b) over the MAG corpus. We also provide the mapping from MAG paper IDs into the raw texts of titles and abstracts here. In addition, all papers are also associated with the year that the corresponding paper was published. The task is to predict the 40 subject areas of arXiv CS papers, e.g., cs.AI, cs.LG, and cs.OS.
The vector $\gamma(u)$ for OGBN-ARXIV is obtained by using our directed graph approach (Section 4.1). The first row of spatial vector $\alpha_{1i}(u)$ is $40$-dimensional, and each entry is the count of citing (incoming) papers from the corresponding class. The second row is $\alpha_{1o}(u)$ and is defined similarly, where each entry is the count of cited (outgoing) papers from the corresponding class. For contextual vector $\beta(u)$, follow the landmark approach described in Section 4.2, employing just one landmark. Since the vectors are weighted vector, we use Euclidean distance to determine the distance between the landmark and a given node. Hence, for each node $\gamma(u)$ has $80$-dimensional spatial, and $40$-dimensional contextual vector, which totals $120$-dimensional ClassContrast vector.
OGBN-MAG: The OGBN-MAG dataset is a heterogeneous network composed of a subset of the Microsoft Academic Graph (MAG) (Wang et al. 2020). It contains four types of entities—papers (736,389 nodes), authors (1,134,649 nodes), institutions (8,740 nodes), and fields of study (59,965 nodes)—as well as four types of directed relations connecting two types of entities—an author is “affiliated with” an institution, an author “writes” a paper, a paper “cites” a paper, and a paper “has a topic of” a field of study. Similar to OGBN-ARXIV, each paper is associated with a 128-dimensional word2vec vector, and all the other types of entities are not associated with input node features. Given the heterogeneous OGBN-MAG data, the task is to predict one of 349 venues (conference or journal) of each paper.
@@
Wikipedia Network (CHAMELEON and SQUIRREL): The datasets depict page-page networks focused on specific topics such as chameleons and squirrels. In these networks, nodes represent articles, and edges denote mutual links between them. Node features are derived from several informative nouns found in the corresponding Wikipedia pages. If a feature is present in the feature list, it signifies the occurrence of an informative noun in the text of the Wikipedia article. The objective is to classify the nodes into five categories based on the average monthly traffic of the respective web pages. In the context of the Wikipedia network, each link between articles does not imply a one-way relationship; instead, it signifies a mutual connection between the two articles, making it an undirected graph. Therefore, undirected feature extraction approaches are employed for $\gamma^{0}(u)$.
@@
Appendix E ClassContrast & Structural Balance Theory
In the ClassContrast framework, the concepts of valleys and peaks serve as a metaphor for the dynamic process by which homophilous graph nodes adjust their positions within an embedding space. These adjustments reflect the nodes’ tendency to connect more strongly with others from the same class, thereby descending into the valleys of their community and distancing themselves from the peaks representing other classes. This behavior can be understood through the lens of an energy landscape, where each node’s position represents a state of potential energy: valleys correspond to low-energy, stable states (nodes strongly connected within their community), and peaks correspond to high-energy, unstable states (nodes weakly connected or misaligned with their community).
Figure 5: The network shown is unbalanced and unclusterable due to the presence of node $v_{7}$ and its edges.
This perspective on node movement and energy states in ClassContrast is conceptually aligned with structural balance theory in social networks. Structural balance theory, originally proposed by Heider (Heider 1946) and later extended by Cartwright and Harary (Cartwright and Harary 1956), addresses the stability of social relationships, particularly in networks where edges between nodes can be either positive (representing friendship) or negative (representing enmity). According to the theory, a balanced state is one in which the network minimizes social tension: friends of friends are expected to be friends, and enemies of friends are expected to be enemies. In graph terms, the theory suggests that triads in the graph should follow a pattern where either all three relationships are positive, or two are negative and one is positive. These configurations are considered balanced and are associated with minimal social tension. The network evolves towards configurations that minimize the overall "social stress" or "tension"—a process not unlike nodes in ClassContrast moving towards a stable, low-energy configuration within their embedding space. For example, consider the network shown in Figure 5. Without $v_{7}$ and its connections, the network would achieve balance where enemies of friends are enemies, and enemies of enemies are friends. To reduce conflict, $v_{7}$ must eventually disconnect either from $v_{6}$ or $v_{4}$.
In both ClassContrast and structural balance theory, the underlying goal is to achieve a state that minimizes some form of conflict as follows.
Energy Minimization in ClassContrast: Homophilous nodes adjust their embeddings in the feature space by forming stronger connections with similar nodes (nodes of the same class). This process is similar to descending into valleys in an energy landscape, where the nodes seek low-energy, stable states. The movement of nodes closer to their communities (valleys) is driven by the desire to minimize energy, which in this context represents the dissimilarity between a node and its community.
Tension Minimization in Structural Balance Theory: In social networks, nodes (representing individuals or entities) aim to achieve balance by adjusting their relationships. An imbalanced triad might resolve by either transforming a negative relationship into a positive one (thus turning an enemy into a friend) or by reinforcing negative ties (further distancing enemies). The network thus moves toward a state where all triads are balanced, minimizing overall social tension.
The analogy between structural balance theory and ClassContrast deepens our understanding of the underlying principles governing networks and also opens up new avenues for research applications. Temporal graph learning is one such application. In real-world applications, networks are often dynamic, with relationships and node attributes changing over time. Understanding how nodes adjust their embeddings in response to these changes (similar to how relationships evolve in structural balance theory) could lead to more efficient temporal models.

### v2

@@ A.2 SCNode Homophily Matrices
A.3 Proof of the Theorem and New Homophily Metrics
In Section˜3.4, we have defined homophily and discussed the relation between SCNode vectors and the homophily notion. Here we give details how to generalize this idea to give finer homophily notions for graphs. First, we give the proof of Theorem˜3.7.
Theorem A.1.
For an undirected graph $\mathcal{G}=(\mathcal{V},\mathbb{E})$, let $\vec{\alpha}_{1}(v)$ be the spatial vector defined in Appendix˜A. Let $\widehat{\alpha}_{1}(v)$ be the vector where the entry corresponding to class of $v$ is set to 0. Then,
@@ A.3 Proof of the Theorem and New Homophily Metrics
For an undirected graph $\mathcal{G}=(\mathcal{V},\mathbb{E})$ with node class assignment function $\mathcal{C}:\mathcal{V}\to\{1,2,\dots,N\}$, $\vec{\alpha}_{1}(v)=[a_{1}(v)\ a_{2}(v)\ \dots\ a_{N}(v)]$ where $a_{i}(v)=\#\{u\in\mathcal{N}_{1}(v)\mid\mathcal{C}(u)=i\}$. Let $\mathcal{C}(v)=j_{v}$. Then, by setting $a_{j_{v}}(v)=0$, we get a new vector $\widehat{\alpha}_{1}(v)=[a_{1}(v)\dots\ a_{j_{v}-1}(v)\ 0\ a_{j_{v}+1}(v)\ \dots]$
@@
where the last equality follows by Equation˜1. The proof follows. ∎
This perspective inspires different ways to generalize the homophily concept by using SCNode vectors. Notice that in SCNode matrices above, we employed a classwise grouping to measure class interactions. If we do not use any grouping for the nodes, we get natural generalizations of existing homophily ratios. First, we define higher homophily by using the ratio of the number of nodes in the $2$-neighborhood $\mathcal{N}_{2}(v)$ with the same class to $|\mathcal{N}_{2}(v)|$.
Definition A.2 (Higher Homophily).
Given $\mathcal{G}=(\mathcal{V},\mathbb{E})$ with $\mathcal{C}:\mathcal{V}\to\{1,2,\dots,N\}$ representing node classes. Let $\eta_{2}(v)$ be the number of nodes in $\mathcal{N}_{2}(v)$ in the same class with $v$. Then, homophily ratio of $\mathcal{G}$ is defined as
@@
By applying similar ideas to the proof of Theorem˜3.7, we obtain the following result.
Theorem A.3.
For an undirected graph $\mathcal{G}=(\mathcal{V},\mathbb{E})$, let $\vec{\alpha}_{2}(v)$ be the spatial vector defined above. Let $\widehat{\alpha}_{2}(v)$ be the vector where the entry corresponding to the class of $v$ is set to 0. Then,
@@
Definition A.4 (Contextual Homophily).
Given $\mathcal{G}=(\mathcal{V},\mathbb{E})$ with $\mathcal{C}:\mathcal{V}\to\{1,2,\dots,N\}$ representing node classes. Let $\vec{\beta}(v)$ be the contextual vector as defined in Section˜3.2. Let $\widehat{\beta}(v)$ be the vector where the entry corresponding to the class of $v$ is set to 0.Then, define the contextual homophily of $\mathcal{G}$ as
@@
Similar homophily metrics, as introduced in Luan et al. (2023); Lim et al. (2021b); Zhu et al. (2020); Jin et al. (2022), were employed to investigate the concept of heterophily in graph representation learning. In particular, our SCNode features inherently encompass this information, enabling the ML classifier to adapt accordingly. This explains the outstanding performance of SCNode in both homophilic and heterophilic settings for node classification and link prediction.
Remark A.5 (SCNode and Underreaching).
Underreaching arises because a depth-$K$ message-passing GNN is $K$-local: its output at node $u$ depends only on the $K$-hop ego-network and features within it, so nodes that are $K$-hop isomorphic remain indistinguishable. SCNode breaks this locality at the input level before message passing.
From a spectral perspective, spatial coordinates $\alpha_{k}(u)$ are class-distribution statistics over $K$-hop neighborhoods and can be expressed as $\alpha_{k}=P_{k}(A)Y$, where $P_{k}$ is a degree-$K$ polynomial in the normalized adjacency $A$ and $Y$ encodes class indicators on labeled nodes. Thus, $\alpha_{k}$ lies within the span of polynomial filters in $A$. In contrast, contextual coordinates $\beta(u)=[d(X_{u},\xi_{1}),\ldots,d(X_{u},\xi_{N})]$ are distances to class landmarks $\{\xi_{j}\}$ in attribute space and are independent of $A$. Concatenating $\alpha_{k}(u)\,\|\,\beta(u)]$ therefore injects a nonlocal, zero-hop channel that cannot be reproduced by any finite-degree polynomial in $A$, strictly enlarging the representable function class.
From an expressivity perspective, two nodes that are $K$-hop isomorphic receive identical inputs under any $K$-local MPGNN. With contextual embeddings $\beta(u)$, such nodes can still be distinguished if they occupy different positions relative to class landmarks in attribute space, allowing even a shallow classifier on $\alpha_{k}\,\|\,\beta]$ to separate them.
Empirically, Table 7 supports this analysis: contextual-only embeddings succeed in heterophilic graphs where local neighborhoods are misleading, while spatial-only embeddings are strongest on homophilic graphs. Their combination consistently achieves the best results, aligning with the theoretical claim that $\beta$ supplies global class information inaccessible to local filters, while $\alpha$ exploits reliable neighborhood structure.
Appendix B Generalizations of SCNode
B.1 Inductive and Transductive Settings
@@
To clarify the differences between the inductive and transductive settings in graph-based learning, consider a given graph $\mathcal{G}$, with datasets $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$. In the inductive setting, all test nodes and their connected edges are removed to create the training subgraph $\mathcal{G}_{\text{train}}$. The model is trained exclusively on $\mathcal{G}_{\text{train}}$ and only gains access to the complete graph $\mathcal{G}$ during the testing stage. Conversely, in the transductive learning approach, the model has access to the entire graph $\mathcal{G}$ during training; however, the labels of the test nodes remain hidden throughout this process. It is noteworthy that any dataset configured for transductive learning can be adapted to the inductive setting by excluding test nodes and their connecting edges during the training phase. However, converting from the inductive to the transductive setting is not generally feasible. For further details, see the references Ciano et al. (2021); Arnold et al. (2007).
B.2 SCNode for Transductive Setting and Iterated Predictions
So far, we outlined our SCNode vectors for simplicity in inductive setting. To adapt to the transductive setting, we make adjustments without altering contextual vectors. In transductive setting, test node labels are hidden during training, but connection information to training nodes is available. We introduce a new "unknown" class $\mathcal{C}_{u}$ for test nodes with unknown labels, considering them as neighbors to training nodes. Each node $u\in\mathcal{V}$ is represented by a $N+1$ dimensional vector $\vec{\alpha}^{0}_{1}(u)=[a^{0}_{10}\ \ a^{0}_{11}\ \ a^{0}_{12}\ \ \dots\ \ a^{0}_{1N}]$, where $a^{0}_{10}$ is the count of neighboring test nodes (unknown labels) of node $u$, and $a^{0}_{1j}$ is the count of neighboring training nodes in class $\mathcal{C}_{j}$ for $1\leq j\leq N$. Similar representations are defined for $\vec{\alpha}_{1i}(u)$, $\vec{\alpha}_{1o}(u)$, and $\vec{\alpha}_{2}(u)$. The superscript 0 indicates no iterations have occurred yet.
Iterated Predictions: Recall that the ultimate goal in the node classification problem is to predict the labels of new (test) nodes. After we obtain all SCNode vectors for training and test nodes above, we let our ML model make a prediction for each test node $v\in\mathcal{V}_{test}$. Let $\mathcal{P}_{0}:\mathcal{V}_{test}\to\{\mathcal{C}_{1},\dots,\mathcal{C}_{N}\}$ be our predictions. Hence, we have a label for each node in our graph $\mathcal{G}$.
@@ B.2 SCNode for Transductive Setting and Iterated Predictions
B.3 Spatial Embeddings on Directed, Weighted Graphs
@@
where $a^{i}_{kj}$ is the count of k-hop neighbors incoming to $u$ belonging to class $\mathcal{C}_{j}$ while $a^{o}_{kj}$ is the count of neighbors outgoing from $u$ belonging to class $\mathcal{C}_{j}$ (See Figure˜2).
@@
Table 11: The dimension of the SCNode embedding used for each dataset in our model
@@
CORA CITESEER PUBMED OGBN-ARXIV OGBN-MAG TEXAS CORNELL WISCONSIN CHAMELEON
Dimension 42 36 112 120 1,524 30 30 30 20
B.4 Further Ablation Studies
In this part, we present further ablation studies to evaluate the sensitivity of SCNode to neighborhood size in spatial embeddings and to similarity metrics in contextual embeddings. Recall that the two embeddings serve distinct roles: (1) Spatial embeddings are derived from neighborhood label distributions up to a fixed hop size and do not involve landmarks; (2) Contextual embeddings are constructed using class landmarks in attribute space, with similarity metrics selected according to feature type (e.g., Jaccard for sparse binary features such as Cora and Citeseer, Euclidean for real-valued features such as PubMed).
Spatial Embeddings: Effect of Neighborhood Size.
We compared spatial embeddings derived from 1-hop and 2-hop neighborhoods. Results show that 2-hop neighborhoods yield modest accuracy improvements across datasets, though at a higher computational cost. For example, on PubMed the accuracy increases from $77.59\pm 0.61$ with 1-hop to $82.28\pm 0.09$ with 2-hop, while runtime grows from 63s to 232s. Based on this trade-off, we adopt 2-hop neighborhoods in the main experiments as a balanced choice. The detailed results are reported in Table 12.
Contextual Embeddings: Effect of Similarity Metrics.
We further evaluated contextual embeddings using Jaccard, cosine, and Euclidean metrics. On binary datasets (Cora, Citeseer), Jaccard provides the best performance, while cosine achieves similar results but at higher computational cost. Euclidean performs substantially worse in this setting, as expected for sparse binary vectors. On real-valued datasets (PubMed), both Euclidean and cosine produce strong results. These findings confirm that the effectiveness of contextual embeddings depends on aligning the similarity metric with the feature type. Full results are provided in Table 13.
These ablations demonstrate that SCNode is robust to landmark quantity and neighborhood size, and that contextual similarity metrics should be selected according to feature type to achieve the best balance between accuracy and efficiency.
Table 12: Effect of Neighborhood Size. Performances of spatial-only vectors used by different k-hop neighborhoods.
Accuracy Time
Ngbd Size Cora Citeseer PubMed Cora Citeseer PubMed
1-hop $83.96\pm 1.35$ $1.28\pm 1.84 77$$7.59\pm 0.61$ 1 s 1 s 63 s
2-hop $84.61\pm 1.20$ $71.82\pm 1.95$ $82.28\pm 0.09$ 3 s 3 s 232 s
Table 13: Effect of Metric Choice. Performances of contextual-only vectors obtained by different similarity metrics.
Accuracy Time
Metric Cora Citeseer PubMed Cora Citeseer PubMed
Jaccard $73.98\pm 2.56$ $67.42\pm 0.85$ n/a 2 s 7 s n/a
Cosine $73.50\pm 2.48$ $67.09\pm 1.80$ $85.62\pm 0.41$ 5 s 15 s 17 s
Euclidean $62.60\pm 2.65$ $60.89\pm 1.75$ $85.92\pm 0.52$ 2 s 6 s 17 s
Appendix C Details of SCNode Embeddings
@@
In all datasets, we basically used the same method to obtain our vectors, however, when the graph type (directed, undirected), or node attribute vector format varies, our methodology naturally adapts the corresponding setting as detailed below. Note that details of OGBN datasets can be found at Hu et al. (2020) and at Stanford’s Open Graph Benchmark site  [^1] https://ogb.stanford.edu/docs/nodeprop/.
@@
CORA: The CORA dataset is a directed graph of scientific publications classified into one of the $7$ classes. Each node $u$ represents a paper and comes with a binary ($0/1$) vector $X_{u}$ of length 1433 indicating the presence/absence of the corresponding word in the paper from a dictionary of 1433 unique words. The task is to predict the subject area of the paper, e.g. Neural Networks, Theory, Case-Based. The directed graph approach is used to extract the attribute vector, $\gamma^{0}(u)$, from the CORA dataset. Recall that this is transductive setting, and the first row is an $8$ dimensional spatial vector $\alpha^{0}_{1i}(u)$ (Section˜3.1) where the first $7$ entries represent the count of citing (incoming) papers from the corresponding class and the $8^{th}$ entry is the count of unknown citing paper $(v_{i}\in D_{test})$ (Section˜B.2). The second row $\alpha^{0}_{1o}(u)$ is defined similarly using the count of cited (outgoing) papers of the corresponding class. The third and fourth row of the spatial feature vector is obtained similarly using the second neighborhood information of each node for citing and cited paper respectively. For the first iteration $\gamma^{1}(u)$, the same setup is followed ignoring the $8^{th}$ entry of each row because there is no unknown class now.
@@
For the contextual vector $\vec{\beta}_{1}(u)$, we follow the landmark approach described in Section˜3.2. We define the first landmark vector $\xi^{1}_{j}$ which is a binary word vector of length 1433 such that the entry is $1$ if the corresponding word is present in any of binary vectors $\mathcal{X}_{u}$ belonging to class $\mathcal{C}_{j}$, and the entry is 0 otherwise. Then, each entry $b_{1j}$ of $\vec{\beta}_{1}(u)$ is the count of common words between $\mathcal{X}_{u}$ and $\xi^{1}_{j}$ for each class, which produces a $7$ dimensional contextual vector. Similarly, for the contextual vector $\vec{\beta}_{2}(u)$, we use a more selective landmark vector $\xi^{2}_{j}$ is defined as a binary word vector of length 1433 indicating the presence/absence of the corresponding word in at least $10\%$ nodes in the class $\mathcal{C}_{j}$. Hence, the initial vector $\gamma^{0}(u)$ is $46$ dimensional ($32$ spatial, $14$ contextual), in the next iterations, $\gamma^{1}(u),\gamma^{2}(u)$ are both $42$-dimensional ($28$ spatial, $14$ contextual).
@@
OGBN-ARXIV: The OGBN-ARXIV dataset is a directed graph, representing the citation network between all Computer Science (CS) arXiv papers indexed by MAG Wang et al. (2020). Each node is an arXiv paper and each directed edge indicates that one paper cites another one. Each paper comes with a 128-dimensional vector obtained by averaging the embeddings of words in its title and abstract. The embeddings of individual words are computed by running the skip-gram model Mikolov et al. (2013) over the MAG corpus. We also provide the mapping from MAG paper IDs into the raw texts of titles and abstracts here. In addition, all papers are also associated with the year that the corresponding paper was published. The task is to predict the 40 subject areas of arXiv CS papers, e.g., cs.AI, cs.LG, and cs.OS.
The vector $\gamma(u)$ for OGBN-ARXIV is obtained by using our directed graph approach (Section˜3.1). The first row of spatial vector $\alpha_{1i}(u)$ is $40$-dimensional, and each entry is the count of citing (incoming) papers from the corresponding class. The second row is $\alpha_{1o}(u)$ and is defined similarly, where each entry is the count of cited (outgoing) papers from the corresponding class. For contextual vector $\beta(u)$, follow the landmark approach described in Section˜3.2, employing just one landmark. Since the vectors are weighted vector, we use Euclidean distance to determine the distance between the landmark and a given node. Hence, for each node $\gamma(u)$ has $80$-dimensional spatial, and $40$-dimensional contextual vector, which totals $120$-dimensional SCNode vector.
OGBN-MAG: The OGBN-MAG dataset is a heterogeneous network composed of a subset of the Microsoft Academic Graph (MAG) Wang et al. (2020). It contains four types of entities—papers (736,389 nodes), authors (1,134,649 nodes), institutions (8,740 nodes), and fields of study (59,965 nodes)—as well as four types of directed relations connecting two types of entities—an author is “affiliated with” an institution, an author “writes” a paper, a paper “cites” a paper, and a paper “has a topic of” a field of study. Similar to OGBN-ARXIV, each paper is associated with a 128-dimensional word2vec vector, and all the other types of entities are not associated with input node features. Given the heterogeneous OGBN-MAG data, the task is to predict one of 349 venues (conference or journal) of each paper.
@@
Wikipedia Network (CHAMELEON): The dataset depict page-page networks focused on specific topics such as chameleons. In these networks, nodes represent articles, and edges denote mutual links between them. Node features are derived from several informative nouns found in the corresponding Wikipedia pages. If a feature is present in the feature list, it signifies the occurrence of an informative noun in the text of the Wikipedia article. The objective is to classify the nodes into five categories based on the average monthly traffic of the respective web pages. In the context of the Wikipedia network, each link between articles does not imply a one-way relationship; instead, it signifies a mutual connection between the two articles, making it an undirected graph. Therefore, undirected feature extraction approaches are employed for $\gamma^{0}(u)$.